OpenStudy (anonymous):

if f(x)=2x+3, then what is the value of the following: f(f(x)) i need help with the steps to get the answer and the answer itself

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

work from the inside out

OpenStudy (misty1212):

\[f(f(x))=f(2x+3)\] then replace \(x\) by \(2x+3\) using parentheses and get \[f(2x+3)=2(2x+3)+3\]

OpenStudy (misty1212):

you can clean it up with some algebra

OpenStudy (anonymous):

thank you!!!